In a ?ABC, if cos A / a = cos B / b = cos C / c and the side a = 2, then area of the triangle is

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
cos A / a = cos B / b = cos C / c
cos A / (2R sin A) = cos B / (2R sin B) = cos C / (2R sin C)
cot A = cot B = cot C
A = B = C
?ABC is equilateral.
Area =Formula for area of an equilateral triangle=root 3
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi